OpenStudy (anonymous):

Simplify (4-3(√2))^2

OpenStudy (anonymous):

\[(4-3\left( \sqrt{2} \right)^2\] 4-3 becomes 1, so you end up with \[\left( \sqrt{2} \right)^2\] and the sqroot cancels with the power, so your answer would be 2

OpenStudy (isaiah.feynman):

Umm not quite..

OpenStudy (isaiah.feynman):

This is an expansion problem.

OpenStudy (isaiah.feynman):

\[(a-b)^{2} = a^{2}-2ab+b^{2}\]

OpenStudy (jim766):

\[(4-3\sqrt{2})(4-3\sqrt{2})\] foil it

OpenStudy (jim766):

\[16 -12\sqrt{2}-12\sqrt{2}+9(2)\]

OpenStudy (anonymous):

so whose correct? and thanks for all who is trying to help

OpenStudy (isaiah.feynman):

Both ways are the same.

OpenStudy (jim766):

mine is not done...need to combine like terms..

OpenStudy (anonymous):

and I dont know if I asked it correctly but its (4-3\[\sqrt{2}\])\[^{2}\]

OpenStudy (anonymous):

dang it

OpenStudy (anonymous):

is 16−122√−122√+9(2) as simple as it can get?

OpenStudy (isaiah.feynman):

Nope.

OpenStudy (isaiah.feynman):

Add the 12s and multiply 2 and 9.

OpenStudy (anonymous):

typed it in my calculator and got 0.0588 which is correct

OpenStudy (isaiah.feynman):

Good. You might wanna do whole numbers next time instead of decimals.

OpenStudy (anonymous):

16√2√2+9(2)?

OpenStudy (anonymous):

16*1+18?

OpenStudy (anonymous):

i feel that is very wrong

OpenStudy (isaiah.feynman):

Yes it is. The simplified form is \[16-24\sqrt{2}+18 = 34-24\sqrt{2}\]

OpenStudy (anonymous):

thank you